OpenStudy (moonlitfate):

A car decelerates from a speed of 64 km/hr to rest in 3.9 seconds. What is the magnitude of the acceleration in m/s^2 to one decimal place?

OpenStudy (abhisar):

Hello @MoonlitFate !

OpenStudy (moonlitfate):

Hello @Abhisar . :D

OpenStudy (abhisar):

Do you know about equation of motions ?

OpenStudy (moonlitfate):

I don't think so?

OpenStudy (abhisar):

Is this equation familiar to you ? \(\sf s=ut+1/2at^2\)

OpenStudy (moonlitfate):

No. :/

OpenStudy (abhisar):

aww..

OpenStudy (moonlitfate):

I'm guessing my lack of knowledge is just going to make this more difficult. :( But, I actually want to learn...

OpenStudy (abhisar):

It's ok everybody at some point lacks knowledge....and the good part is we can always learn ;)

OpenStudy (moonlitfate):

True. I wish I would have studied over the summer instead of jumping in... to Physics and Calculus II. SO difficult...

OpenStudy (abhisar):

So there are three equations of motion when the acceleration is constant, \(\sf v=u+at\\s=ut+1/2at^2\\v^2=u^2+2as\)

OpenStudy (abhisar):

Here, u=initial velocity, v=final velocity, a=accceleration, s=distance covered and t=time taken

OpenStudy (moonlitfate):

Okay. I'm just writing this down as you go. Sill here.

OpenStudy (abhisar):

ok that's cool. For solving our question at the top, we are gonna use the second equation but let's first understand the question...u'll see how easy that question is !

OpenStudy (abhisar):

A car decelerates from a speed of 64 km/hr to rest in 3.9 seconds. What is the magnitude of the acceleration in m/s^2 to one decimal place? So, the initial velocity of the car is 64km/hr or 17.7 m/s. The decelerates finally to rest so the final velocity will be 0. Time taken for this is 3.9 seconds. ok ?

OpenStudy (moonlitfate):

Yes, get going. Copying this done, and re-reading to make sure I understand. :)

OpenStudy (abhisar):

ok, now we simply plug in the value into the first equation of motion. v=0 u=17.7 t=3.9 => 0=17.7 +a*3.9 Solving it for a, we get -17.7/3.9=a = -4.38 m/s^2 \(\therefore\) The acceleration is \(\sf -4.38m/s^2\) or the deceleration is \(\sf 4.38 m/s^2\)

OpenStudy (moonlitfate):

Oh, that makes sense since. :)

OpenStudy (abhisar):

are you sure ?

OpenStudy (moonlitfate):

Yeah, can we do some more though? :)

OpenStudy (abhisar):

yeah sure ! and To know more about the equations, you can visit http://hyperphysics.phy-astr.gsu.edu/hbase/mot.html

OpenStudy (moonlitfate):

Thank you! :) Can I type up the next one?

OpenStudy (abhisar):

Yeah sure ! Type the next one as a new post :)

OpenStudy (moonlitfate):

Alright.